OpenStudy (anonymous):

If sin(1/2x)=sqrt((x-1)/(2x)), what is tanx?

OpenStudy (anonymous):

\[\sin(\frac{ 1 }{ 2 }\theta)=\sqrt{\frac{ x-1 }{ 2x }}\]solve for \[\tan \theta=? \]in terms of x

OpenStudy (raden):

try represented it in figure :  u have to determind the length of adj of angle x/2